what is trignometric? how many identities are there?

Grade:9

1 Answers

Rinkoo Gupta
9 years ago
Trigonometry is the study of relations between angles and sides of a triangle. The basic trigonometrical identities are:
(1) sin2x+cos2x=1
(2) sec2x=1+tan2x
(3)cosec2x=1+cot2x

Thanks & Regards
Rinkoo Gupta
AskIITians Faculty